The Saturn V was the most powerful rocket ever built and was central to the success of the Apollo program in the 1960s. The rocket used liquid oxygen (O2) as a component of its propellant systems in all three stages.
Consider a liquid oxygen (O2) tank within the rocket of volume 443968 L. How many molecules of O2 would be present in this tank? The density of liquid oxygen is 1.14 kg/L.

Express your answer to two significant figures.
1
Expert's answer
2020-05-25T14:12:56-0400

Applying formula

"density=\frac{mass}{volume }"

"mass =1.14\times443968=506123.52 kg""=506123520grams"


Let's calculate the moles of oxygen

no. of moles"=\frac{given\ mass }{molar\ mass }=\frac{506123520}{32}=15816360=1.58\times10^7 moles"


1 mole contains "6.023\times10^{23}" molecules

"1.58\times10^7" mole will contain "1.58\times10^7\times6.023\times10^{23} molecules =9.51\times10^{30}" molecules